# Model averaging with mixed criteria for estimating high quantiles of extreme values: Application to heavy rainfall

### v1

@@ Abstract
Accurately estimating high quantiles beyond the largest observed value is crucial in risk assessment and devising effective adaptation strategies to prevent a greater disaster. The generalized extreme value distribution is widely used for this purpose, with L-moment estimation (LME) and maximum likelihood estimation (MLE) being the primary methods. However, estimating high quantiles with a small sample size becomes challenging when the upper endpoint is unbounded, or equivalently, when there are larger uncertainties involved in extrapolation. This study introduces an improved approach using a model averaging (MA) technique. The proposed method combines MLE and LME to construct candidate submodels and assign weights effectively. The properties of the proposed approach are evaluated through Monte Carlo simulations and an application to maximum daily rainfall data in Korea. Additionally, theoretical considerations are provided, including asymptotic variance with random weights. A surrogate model of MA estimation is also developed and applied for further analysis.
@@ 1 Introduction
The generalized extreme value (GEV) distribution (GEVD) is extensively used to model extremes in natural phenomena and human activities, such as hydrological events and incidents in insurance and financial markets (Coles 2001; Katz et al. 2002; Reiss & Thomas 2007). The primary reason for applying the GEVD is that it serves as a large-sample approximation to the distribution of sample maximum, regardless of the population from which the data are drawn. The cumulative distribution function (CDF) of the GEVD is given by (Hosking & Wallis 1997):
@@
when $01-\xi(x-\mu)/\sigma>$ 0 and $\sigma>0$, where $\mu,\;\sigma$, and $\xi$ represent the location, scale, and shape parameters, respectively. The case where $\xi=0$ in (1) corresponds to the Gumbel distribution. Notably, the sign of $\xi$ in (1) differs from that used in Coles (2001) to align with the notation in Hosking & Wallis (1997). The tail behavior of the GEV distribution varies depending on the sign of $\xi$: it exhibits a heavy tail when $\xi<0$, an exponential tail as $\xi\rightarrow 0$, and light or bounded when $\xi>0$.
@@
Accurately estimating high upper quantiles is more critical than estimating the parameters of extreme value distributions, particularly in risk assessment and hydraulic structure design (Naghettini 2017; Allouche et al. 2023). For example, the 1% chance flood is computed at the 0.99 quantile of an extreme value distribution which is actually used by the US Federal Emergency Management Agency when developing its regulatory policy for floodplains (Grego & Yates 2024). When the upper endpoint is unbounded (or $\xi<0$), estimating high quantiles of the GEVD with a small sample size becomes challenging because of larger uncertainties involved in model extrapolation. For instance, when $\xi=-0.35$ and the true $0.99$ quantile is 443 (with $\mu=100$ and $\sigma=30$), the root mean squared errors (RMSEs) of the MLE and LME of $0.99$ quantile with $n=60$ are approximately 148 and 120, respectively. These RMSEs represent 33% and 27% of the true quantile, which is concerned as a substantial uncertainty.
@@
In the context of the GEVD for block maxima data, several approaches have been proposed to address this issue, including a mixed estimation of the MLE and LME (Morrison & Smith 2002; Ailliot et al. 2011) and the application of a penalty function to the GEVD parameters (Coles & Dixon 1999; Bücher et al. 2021). The former method effectively and robustly estimates the parameters while reducing RMSE, whereas the latter can outperform MLE in certain situations depending on the choice of penalty function. Lekina et al. (2014) and Fabiola et al. (2021) introduced a semi-parametric and nonparametric weighted estimations of high quantiles, respectively, offering more flexibility and suitability than parametric methods. Daouia et al. (2024) presented general weighted pooled estimators of extreme quantiles calculated through a nonstandard geometric averaging scheme. These weighted estimation methods perform well for large sample sizes. Karvanen (2006) considered an estimation of quantile mixtures via L-moments and trimmed L-moments. Grego et al. (2015) used finite mixture models to accommodate annual peak streamflow. Evin et al. (2011) and Totaro et al. (2024) explored using two-component mixtures of Gumbel distributions. Our study builds upon this framework by introducing a modified approach that employs mixtures of narrow extreme value submodels.
@@
In this study, we propose a new estimation method to improve the accuracy of high quantile estimation in the GEVD using model averaging (MA). The MA combines the strengths of multiple candidate models by assigning greater weights to superior models (Liu et al. 2016; Fletcher 2018). Unlike traditional model selection, which discards all but the best model, MA explicitly accounts for model uncertainty by integrating information from competing statistical models. MA reduces the risk of model misspecification and makes predictions less sensitive to model assumptions. By incorporating model uncertainty, MA produces more accurate estimates and reliable confidence intervals than model selection from frequentist and Bayesian perspectives (Burkland et al. 1999; Hoeting et al. 1999; Claeskens & Hjort 2008). Le and Clarke (2022) showed that MA is asymptotically better than model selection for prediction.
@@
However, the MA method has some limitations, such as computational cost. The slowest candidate model indicates the overall computational speed of MA. Additionally, MA does not consistently outperform model selection. Model selection is preferred when the error variance is small and identifying the best model is relatively straightforward. Conversely, MA performs well when the error variance is large (Liu et al. 2016, 2023). Given the high error variance in estimating high quantiles of the GEVD under a negative shape parameter, MA may outperform some existing methods.
@@
Several Bayesian MA approaches for extreme values have been developed based on the likelihood function and posterior probability (Raftery et al. 2005; Sabourin et al. 2013; Vettori et al. 2020). These approaches typically estimate the parameters of candidate models using MLE or Bayes estimators and assign weights based on marginal likelihood. In frequentist MA, candidate model parameters are estimated using the least squares method or MLE, and weights are determined based on information criteria, cross-validation, or an unbiased risk estimator (Claeskens & Hjort 2008; Wang et al. 2009).
@@
This study applies different criteria for estimation and weighting in MA, primarily utilizing L-moments. This mixed application of criteria is a key aspect of the proposed method, which was motivated by Morrison & Smith (2002) and Ailliot et al. (2011). We hypothesize that the mixed-criteria approach—where candidate model parameters are estimated using MLE and weights are assigned based on L-moments, or vice versa—may outperform the conventional single-criterion method, which estimates parameters via MLE and assigns weights based on the likelihood function. We anticipate that this mixed-criteria approach will improve the estimation of high quantiles of extreme values. To our knowledge, L-moment-based MA methods have been scarcely studied, despite the popularity of non-MA approaches based on L-moments in climatology and hydrology (Hosking & Wallis 1997; Naghettini 2007; Kjeldsen et al. 2017; Hong et al. 2022; Strong et al. 2025, for example among many). We consider two methods: (1) estimating candidate model parameters using MLE and calculating weights using the generalized L-moments distance, and (2) estimating candidate model parameters using LME and calculating weights using the smooth Akaike information criterion. For comparison, we also evaluate other methods, including mixed estimation (Morrison & Smith 2002; Ailliot et al. 2011) and a penalized likelihood method (Coles & Dixon 1999).
@@
The rest of this paper is structured as follows: Section 2 details the proposed MA methods. Theoretical considerations on the MA estimator and uncertainty calculation under random weights are presented in Section 3. Section 4 provides a simulation study. Section 5 presents an application using maximum rainfall data from Hae-nam, Korea. Discussion and conclusion are provided in Sections 6 and 7. Supplementary Material includes additional details such as formulas, tables, and figures. The R code for our implementation is available on GitHub: https://github.com/yire-shin/MA-gev.git.
@@ 2 Model averaging methods
The MA methodology consists of three steps: (1) constructing candidate models, (2) estimating model parameters, and (3) calculating weights for each candidate model (Liu et al. 2023; Salaki et al. 2024; Hao et al. 2024). The primary objective of MA in this study is to improve the estimation of high quantiles of extreme values rather than to focus on estimating the parameters of the GEVD. The $1/p$ return level, denoted as $r_{1/p}$, represents the $1-p$ quantile of the distribution (Coles 2001). The “T-year return level” is commonly used for annual extreme values, where $T=1/p$. The MLE and LME of the three GEVD parameters are defined as follows:
@@ 2.1 Construction of candidate models
A concrete application of MA often requires an initial selection of several candidate models. These candidate models may differ entirely or only in certain components. Among three parameters in GEVD, estimating the shape parameter $\xi$ is more challenging than others. Additionally, $\xi$ is dimensionless, and the behavior of GEVD is more sensitive to $\xi$ than the other two parameters. Therefore, the proposed method starts with the shape parameter $\xi$. We consider two approaches for initializing our MA algorithm: the MLE of $\xi$ and the LME of $\xi$.
In the first approach, based on $\hat{\xi}_{M}$, we construct a $100\times(1-\alpha)$ confidence interval for $\xi$ using the profile likelihood (Coles 2001). For instance, setting $\alpha=0.05$ provides a 95% confidence interval $(C_{L},\;C_{U})$. We select $K$ values of $\xi$ proportionally distributed according to the profile likelihood of $\xi$ within this interval from $C_{L}$ to $C_{U}$, since values near the lower and upper bounds are less probable than those in the interior. In the second approach, based on $\hat{\xi}_{L}$, we use a nonparametric bootstrap to construct a $100\times(1-\alpha)$ confidence interval for $\xi$. Subsequently, we select $K$ values of $\xi$ proportionally (with equal probability) from the bootstrap distribution of $\hat{\xi}_{L}$ within this confidence interval. This study considers two values of $\alpha$: $0.05$ and $0.10$. The upper panel of Figure 1 illustrates how the proposed method selects ten values of $\xi_{k}$s using a 95% confidence interval of $\xi$ based on the profile likelihood function.
For each selected $\xi_{k}$, we estimate $\mu_{k}$ and $\sigma_{k}$ using either the MLE or LME, keeping $\xi_{k}$ fixed. These estimators are denoted as $\hat{\mu}_{k\text{M}},\;\hat{\sigma}_{k\text{M}}$, and $\hat{\mu}_{k\text{L}},\;\hat{\sigma}_{k\text{L}}$, respectively. Estimating two other parameters with $\xi_{k}$ fixed is computationally simpler and yields lower standard errors (SE) than estimating all three GEVD parameters. When $\xi$ is known, the LMEs of $\mu_{k}$ and $\sigma_{k}$ are given by:
@@
$\hat{\mu_{k}}=l_{1}-\hat{\sigma_{k}}\ \{1-\Gamma(1+\xi_{k})\}/\xi_{k},$
where $l_{1}$ and $l_{2}$ are the first and second sample L-moments computed from the data. As a result, we obtained $K$ GEV candidate models, where each is the two-parameter GEVD. From each candidate model, we compute the T-year return level ($\hat{r}_{T,k}$) using the estimated parameters $(\hat{\mu}_{k\text{M}},\;\hat{\sigma}_{k\text{M}},\;\xi_{k})$ or $(\hat{\mu}_{k\text{L}},\;\hat{\sigma}_{k\text{L}},\;\xi_{k})$. Finally, we compute the MA T-year return levels by averaging these T-year return levels with weights $w_{k}$, which are nonnegative and summed to one.
@@ 2.2 Weighting scheme based on L-moments distance
When the parameters in each candidate model are estimated using the MLE, reusing the likelihood value to compute weights—such as through the smooth Akaike information criterion (AIC) (Claeskens & Hjort 2008; Fletcher 2018)—may not necessarily improve estimation, in our view. As demonstrated later in this study, conventional model averaging that employs MLE for parameter estimation and smooth AIC for weight assignment does not perform well, at least in our simulation study. Therefore, we explored alternative criteria to construct more informative weights.
For a three-parameter distribution, the L-moment distance for each candidate model $M_{k}$ is defined as follows, subject to the constraint that $\xi>-1$ for the GEVD:
@@
Since the LME is obtained by solving equations that ensure (4) equals zero, we aim to assign higher weights to submodels with smaller L-moment distances. This weighting scheme is preferable because L-moments characterize the distribution differently from the likelihood function.
@@
where $d_{k}^{T}=(l_{1}-\lambda_{1,k},\;l_{2}-\lambda_{2,k},\;l_{3}-\lambda_{3,k})$, and $V_{k}$ is a $3\times 3$ covariance matrix of $d_{k}$. A similar GLD approach has been explored in previous studies, including Kjeldsen & Prosdocimi (2015) for regional frequency analysis as well as Alvarez et al. (2022).
To compute $V_{k}$ for each model, one can use a parametric bootstrap based on the GEVD with parameter estimates $(\hat{\mu}_{k\text{M}},\;\hat{\sigma}_{k\text{M}},\;\xi_{k})$ since an explicit form of $V_{k}$ is unavailable. However, this approach is computationally intensive. To mitigate this issue, we approximated $V_{k}$ using the covariance matrix of the sample L-moments (Elamir & Seheult 2004), treating $\lambda_{j,k}$ as constants. For this computation, we employed the ‘lmomco’ package (Asquith 2014) in R. However, in some cases, this method resulted in a non-positive definite covariance matrix. In such instances, we computed the covariance matrix using a nonparametric bootstrap. Once computed, $V$ was fixed and applied uniformly across all $K$ candidate models, as it no longer depended on $k$.
To assign higher weights to submodels with smaller $\text{GLD}_{k}$ values, we adopted the multivariate normal probability:
@@
$\hat{w}_{k}={\frac{p(\>\underline{x}|M_{k})}{\sum_{k=1}^{K}p(\>\underline{x}|M_{k})}},$ (7)
which ensures that $\sum_{k=1}^{K}w_{k}=1$ and that all $w_{k}$ values remain within the range $[0,1]$. This weight can be interpreted as the probability that model $M_{k}$ is the best among the set of candidate models. Given that sample L-moments generally converge to a multivariate normal distribution as $n\rightarrow\infty$ (Hosking 1990), the function in (6) serves as an approximation of the likelihood function of population L-moments given three sample L-moments.
For a more robust L-moment distance, one can replace $l_{1}$ and $l_{2}$ with the median and interquartile range of the data. While $l_{2}$ is proportional to Gini’s mean difference and thus already serves as a robust scale estimate, $l_{1}$ (the sample mean) is not (Ailliot et al. 2011). To address this, we substitute the first element of $d_{k}$ with the median distance (sample $\text{sample median}-q_{0.5}$) as an alternative to (5). In the following sections, we refer to this alternative weighting scheme as the ‘med’ weight.
The lower panel of Figure 1 illustrates how the proposed method computes the return level estimate using a weighted average of return values from 10 submodels. The 200-year return level obtained via the MA method lies between the estimates derived from the MLE and LME.
Figure 1: Illustration of the selection of $\xi_{k}$s values using the profile likelihood of $\xi$ and the corresponding weights, based on annual maximum daily rainfall data (unit: $mm$) from Hae-nam, Korea. (Top panel): The green curve represents the profile log-likelihood function, while the 10 vertical lines indicate the selected $\xi_{k}$s values. (Bottom panel): The 200-year return level estimates from 10 submodels (vertical dotted lines) are shown along with their corresponding weights (black dots) for the proposed MA method (‘like1’). The return level estimates obtained from the MLE, LME, and MA methods are also displayed for comparison.

### v2

@@ Abstract
Accurately estimating high quantiles beyond the largest observed value is crucial for risk assessment and devising effective adaptation strategies to prevent a greater disaster. The generalized extreme value distribution is widely used for this purpose, with L-moment estimation (LME) and maximum likelihood estimation (MLE) being the primary methods. However, estimating high quantiles with a small sample size becomes challenging when the upper endpoint is unbounded, or equivalently, when there are larger uncertainties involved in extrapolation. This study introduces an improved approach using a model averaging (MA) technique. The proposed method combines MLE and LME to construct candidate submodels and assign weights effectively. The properties of the proposed approach are evaluated through Monte Carlo simulations and an application to maximum daily rainfall data in Korea. In addition, theoretical properties of the MA estimator are examined, including the asymptotic variance with random weights. A surrogate model of MA estimation is also developed and applied for further analysis. Finally, a Bayesian model averaging approach is considered to reduce the estimation bias occurring in the MA methods.
@@ 1 Introduction
The generalized extreme value (GEV) distribution (GEVD) is extensively used to model extremes in natural phenomena and human activities, such as hydrological events and incidents in insurance and financial markets (Coles 2001; Katz et al. 2002; Reiss & Thomas 2007). The primary reason for applying the GEVD is that it serves as a large-sample approximation to the distribution of sample maxima, regardless of the population from which the data are drawn. The cumulative distribution function (CDF) of the GEVD is given by (Hosking & Wallis 1997):
@@
when $1-\xi(x-\mu)/\sigma>0$ and $\sigma>0$, where $\mu,\;\sigma$, and $\xi$ represent the location, scale, and shape parameters, respectively. The case where $\xi=0$ in (1) corresponds to the Gumbel distribution. Notably, the sign of $\xi$ in (1) differs from that used in Coles (2001) to align with the notation in Hosking & Wallis (1997). The tail behavior of the GEV distribution varies depending on the sign of $\xi$: it exhibits a heavy tail when $\xi<0$, an exponential tail as $\xi\rightarrow 0$, and light or bounded when $\xi>0$.
@@
Accurately estimating high upper quantiles is more critical than estimating the parameters of extreme value distributions, particularly in risk assessment and the design of hydraulic structure (Naghettini 2017; Allouche et al. 2023; Zheng et al. 2025). For example, the 1% chance flood is computed at the 0.99 quantile of an extreme value distribution which is actually used by the US Federal Emergency Management Agency when developing its regulatory policy for floodplains (Grego & Yates 2025). When the upper endpoint is unbounded (i.e., $\xi<0$), estimating high quantiles of the GEVD with a small sample size becomes challenging because of larger uncertainties associated with model extrapolation. For instance, when $\xi=-0.35$ and the true $0.99$ quantile is 443 (with $\mu=100$ and $\sigma=30$), the root mean squared errors (RMSEs) of the MLE and LME of the $0.99$ quantile with $n=60$ are approximately 148 and 120, respectively. These RMSEs represent 33% and 27% of the true quantile, which is considered a substantial estimation uncertainty.
@@
In the context of the GEVD for block maxima data, several approaches have been proposed to address this issue, including a mixed estimation of the MLE and LME (Morrison & Smith 2002; Ailliot et al. 2011) and the application of a penalty function to the GEVD parameters (Coles & Dixon 1999; Bücher et al. 2021). The former method estimates the parameters effectively and robustly while reducing RMSE, whereas the latter can outperform MLE in certain situations depending on the choice of penalty function. Lekina et al. (2014) and Fabiola et al. (2021) introduced semi-parametric and nonparametric weighted estimations of high quantiles, respectively, offering more flexibility and suitability than parametric methods. Daouia et al. (2024) presented general weighted pooled estimators of extreme quantiles calculated through a nonstandard geometric averaging scheme. These weighted estimation methods generally perform well for large sample sizes. Karvanen (2006) considered the estimation of quantile mixtures via L-moments and trimmed L-moments. Grego et al. (2015) used finite mixture models to accommodate annual peak streamflow. Evin et al. (2011) and Totaro et al. (2024) explored the use of two-component mixtures of Gumbel distributions. Our study builds upon this framework by introducing a modified approach that employs mixtures of narrow extreme value submodels.
@@
In this study, we propose a new estimation method aimed at improving the accuracy of high quantile estimation in the GEVD using model averaging (MA). The MA combines the strengths of multiple candidate models by assigning greater weights to superior models (Liu et al. 2016; Fletcher 2018). Unlike traditional model selection, which discards all but the best model, MA explicitly accounts for model uncertainty by integrating information from competing statistical models. MA reduces the risk of model misspecification and makes predictions less sensitive to model assumptions. By incorporating model uncertainty, MA produces more accurate estimates and more reliable confidence intervals than model selection from frequentist and Bayesian perspectives (Burkland et al. 1999; Hoeting et al. 1999; Claeskens & Hjort 2008). Le and Clarke (2022) showed that MA is asymptotically better than model selection in terms of predictive performance.
@@
However, the MA method has some limitations, such as computational cost. The slowest candidate model largely determines the overall computational speed of MA. Additionally, MA does not consistently outperform model selection. Model selection is preferred when the error variance is small and identifying the best model is relatively straightforward. Conversely, MA performs well when the error variance is large (Liu et al. 2016, 2023). Given the high error variance in estimating high quantiles of the GEVD under a negative shape parameter, MA is expected to outperform some existing methods.
@@
Several Bayesian MA approaches for extreme values have been developed based on the likelihood function and posterior probability (Raftery et al. 2005; Sabourin et al. 2013; Vettori et al. 2020). These approaches typically estimate the parameters of candidate models using MLE or Bayes estimators and assign weights based on the marginal likelihood. In frequentist MA, candidate model parameters are estimated using the least squares method or MLE, and weights are determined based on information criteria, cross-validation, or an unbiased risk estimator (Claeskens & Hjort 2008; Wang et al. 2009).
@@
This study applies different criteria for estimation and weighting in MA, with a primary focus on L-moments. This mixed application of criteria is a key aspect of the proposed method, which was motivated by Morrison & Smith (2002) and Ailliot et al. (2011). We hypothesize that the mixed-criteria approach—where candidate model parameters are estimated using MLE and weights are assigned based on L-moments, or vice versa—can outperform the conventional single-criterion method, which estimates parameters via MLE and assigns weights based on the likelihood function. We anticipate that this mixed-criteria approach will improve the estimation of high quantiles of extreme values. To our knowledge, L-moment-based MA methods have been rarely studied, despite the popularity of non-MA approaches based on L-moments in climatology and hydrology (Hosking & Wallis 1997; Naghettini 2007; Kjeldsen et al. 2017; Lee et al. 2020; Strong et al. 2025, among others).
We consider two methods: (1) estimating candidate model parameters using MLE and calculating weights using the generalized L-moment distance, and (2) estimating candidate model parameters using LME and calculating weights using the smooth Akaike information criterion. For comparison, we also evaluate other methods, including mixed estimation (Morrison & Smith 2002; Ailliot et al. 2011) and a penalized likelihood method (Coles & Dixon 1999).
@@
The rest of this paper is structured as follows: Section 2 details the proposed MA methods. Theoretical considerations on the MA estimator and uncertainty calculation under random weights are presented in Section 3. Section 4 provides a simulation study. Section 5 presents an application using maximum rainfall data from Hae-nam, Korea. Further development including the selection of the number of submodels and bias correction using Bayesian model averaging are presented in Section 6. Discussion and conclusion are provided in Sections 7 and 8. The Supplementary Material includes additional details such as formulas, tables, and figures. The R code for our implementation is available on GitHub: https://github.com/yire-shin/MA-gev.git.
@@ 2 Model averaging methods
The MA methodology consists of three steps: (1) constructing candidate models, (2) estimating model parameters, and (3) calculating weights for each candidate model (Liu et al. 2023; Salaki et al. 2024; Hao et al. 2024). The primary objective of MA in this study is to improve the estimation of high quantiles of extreme values, rather than to focus on parameter estimation of the GEVD. The $1/p$ return level, denoted as $r_{1/p}$, represents the 1p1-p quantile of the distribution (Coles 2001). The “T-year return level” is commonly used for annual extreme values, where $T=1/p$. The MLE and LME of the three GEVD parameters are defined as follows:
@@ 2.1 Construction of candidate models
A concrete application of MA often requires an initial selection of several candidate models. These candidate models may differ entirely or only in certain components. Among the three parameters in the GEVD, estimating the shape parameter $\xi$ is more challenging than estimating the other two parameters. Additionally, $\xi$ is dimensionless, and the behavior of the GEVD is more sensitive to $\xi$ than to the other two parameters. Therefore, the proposed method starts with the shape parameter $\xi$. We consider two approaches for initializing our MA algorithm: the MLE of $\xi$ and the LME of $\xi$.
In the first approach, based on $\hat{\xi}_{M}$, we construct a $100\times(1-\alpha)$ confidence interval for $\xi$ using the profile likelihood (Coles 2001). For instance, setting $\alpha=0.05$ provides a 95% confidence interval $(C_{L},\;C_{U})$. We select $K$ values of $\xi$ proportionally to the profile likelihood of $\xi$ within this interval from $C_{L}$ to $C_{U}$, since values near the lower and upper bounds are less probable than in the interior of the interval. The upper panel of Figure 1 illustrates how the proposed method selects ten values of $\xi_{k}$ using a 95% confidence interval of $\xi$ based on the profile likelihood function. In the second approach, based on $\hat{\xi}_{L}$, we use a nonparametric bootstrap to construct a $100\times(1-\alpha)$ confidence interval for $\xi$. Subsequently, we select $K$ values of $\xi$ with equal probability from the bootstrap distribution of $\hat{\xi}_{L}$ within this confidence interval. This study considers two values of $\alpha$: $0.05$ and $0.10$.
For each selected $\xi_{k}$, we estimate $\mu_{k}$ and $\sigma_{k}$ using either the MLE or LME, keeping $\xi_{k}$ fixed. These estimators are denoted as $\hat{\mu}_{k\text{M}},\;\hat{\sigma}_{k\text{M}}$, and $\hat{\mu}_{k\text{L}},\;\hat{\sigma}_{k\text{L}}$, respectively. Estimating the two other parameters with $\xi_{k}$ fixed is computationally simpler and yields lower standard errors (SE) than estimating all three GEVD parameters. When $\xi$ is known, the LMEs of $\mu_{k}$ and $\sigma_{k}$ are given by:
@@
[…]
where $l_{1}$ and $l_{2}$ are the first and second sample L-moments computed from the data. As a result, we obtained $K$ GEV candidate models, where each is a two-parameter GEVD. From each candidate model, we compute the T-year return level ($\hat{r}_{T,k}$) using the estimated parameters $(\hat{\mu}_{k\text{M}},\;\hat{\sigma}_{k\text{M}},\;\xi_{k})$ or $(\hat{\mu}_{k\text{L}},\;\hat{\sigma}_{k\text{L}},\;\xi_{k})$. Finally, we compute the MA T-year return levels by averaging these T-year return levels with weights $w_{k}$, which are nonnegative and sum to one.
Figure 1: Illustration of the selection of $\xi_{k}$ values using the profile likelihood of $\xi$ and the corresponding weights, based on annual maximum daily rainfall data (unit: $mm$) from Hae-nam, Korea. (Top panel): The green curve represents the profile log-likelihood function, while the ten vertical lines indicate the selected $\xi_{k}$ values. (Bottom panel): The 200-year return level estimates from ten submodels (vertical dotted lines) are shown along with their corresponding weights (black dots) for the proposed MA method (‘like1’). The return level estimates obtained from the MLE, LME, and MA methods are also displayed for comparison.
@@ 2.2 Weighting scheme based on L-moments distance
When the parameters in each candidate model are estimated using the MLE, reusing the likelihood value to compute weights—such as through the smooth Akaike information criterion (AIC) (Claeskens & Hjort 2008; Fletcher 2018)—may not necessarily improve estimation, in certain settings. As demonstrated later in this study, conventional model averaging that employs MLE for parameter estimation and smooth AIC for weight assignment does not yield satisfactory performance, at least in our simulation study. Therefore, we explored alternative criteria to construct weights that are more informative for high-quantile estimation.
For a three-parameter distribution, the L-moment distance for each candidate model $M_{k}$ is defined as follows, subject to the constraint that $\xi>-1$ in the GEVD:
@@
Since the LME is obtained by solving equations that ensure $LD_{k}=0$, we aim to assign higher weights to submodels with smaller L-moment distances. This weighting scheme is preferable because L-moments characterize the distribution differently from the likelihood function.
@@
where $d_{k}^{T}=(l_{1}-\lambda_{1,k},\;l_{2}-\lambda_{2,k},\;l_{3}-\lambda_{3,k})$, and $V_{k}$ is a $3\times 3$ covariance matrix of $d_{k}$. A similar GLD approach has been explored in previous studies, including Kjeldsen & Prosdocimi (2015) for regional frequency analysis as well as Alvarez et al. (2022), Shin & Park (2024), and Shin et al. (2025a).
To compute $V_{k}$ for each model, one can use a parametric bootstrap based on the GEVD with parameter estimates $(\hat{\mu}_{k\text{M}},\;\hat{\sigma}_{k\text{M}},\;\xi_{k})$ since an explicit form of $V_{k}$ is unavailable. However, this approach is computationally demanding. To mitigate this issue, we approximate $V_{k}$ using the covariance matrix of the sample L-moments (Elamir & Seheult 2004), treating $\lambda_{j,k}$ as constants. For this computation, we employed the ‘lmomco’ package (Asquith 2014) in R. However, in some cases, this method resulted in a non-positive definite covariance matrix. In such instances, we computed the covariance matrix using a nonparametric bootstrap. Once computed, $V$ was fixed and applied uniformly across all $K$ candidate models, as it no longer depended on $k$.
To assign higher weights to submodels with smaller $\text{GLD}_{k}$ values, we adopted the multivariate normal density:
@@
$\hat{w}_{k}={\frac{p(\>\underline{x}|M_{k})}{\sum_{j=1}^{K}p(\>\underline{x}|M_{j})}},$ (7)
which ensures that $\sum_{k=1}^{K}\hat{w}_{k}=1$ and that all $\hat{w}_{k}$ values remain within the range $[0,1]$. This weight can be interpreted as the probability that model $M_{k}$ is the best among the set of candidate models. Given that sample L-moments generally converge to a multivariate normal distribution as $n\rightarrow\infty$ (Hosking 1990), the function in (6) serves as an approximation of the likelihood function of population L-moments given three sample L-moments.
For a more robust L-moment distance, one can replace $l_{1}$ and $l_{2}$ with the sample median and interquartile range of the data. While $l_{2}$ is proportional to Gini’s mean difference and thus already serves as a robust scale estimate, $l_{1}$ (the sample mean) is not (Ailliot et al. 2011). To address this, we substitute the first element of $d_{k}$ with the median-based distance ($\text{sample median}-q_{0.5}$) as an alternative to (5). In the following sections, we refer to this alternative weighting scheme as the ‘med’ weight.
The lower panel of Figure 1 illustrates how the proposed method computes the return level estimate using a weighted average of return values from ten submodels. The 200-year return level obtained via the MA method lies between the estimates derived from the MLE and LME.